OpenStudy (anonymous):

# A 5-foot tall student stands 10 feet away from the lens of a camera. The film of the camera is 3 inches away from the lens. A 5 ft tall student is standing 10 feet away from the lens of a camera; behind the lens of the camera is shown the film of the camera, 3 inches away from the lens. What is the height of the film? Answer 0.2 inches 0.7 inches 1.2 inches 1.5 inches

hero (hero):

Was there a picture included with this?

OpenStudy (anonymous):

yes. one second

hero (hero):

What exactly is the topic of study for your class?

OpenStudy (anonymous):

meaning? Its geometry and that lesson was about similar triangles

hero (hero):

That's what I needed to know

OpenStudy (anonymous):

kk

hero (hero):

Okay, I figured out what it was, but the explanation is kind of whacky

hero (hero):

You're supposed to create a proportion relating the two triangles by setting height of the person to the height of the film equal to the length of the person from the lens to the length of the film from the lens.

OpenStudy (anonymous):

right. I can get there but whats next?

OpenStudy (anonymous):

do I cross multiply?

hero (hero):

Next you cross multiply, then solve for x.

hero (hero):

wait a second....

hero (hero):

We need to convert feet to inches.

OpenStudy (anonymous):

oh lord. I didnt notice that......

hero (hero):

Yeah, I just realized it myself.

OpenStudy (anonymous):

1/40 = 5/x ??

OpenStudy (anonymous):

x = 200.........right?

hero (hero):

So after converting ft to inches, you get 60 for the height of the person 120 the person's distance from the lens So replacing those in the proportion you get: \[\frac{60}{x}=\frac{120}{3}\]

OpenStudy (anonymous):

Just kidding! geez my brain is fried

hero (hero):

\[\frac{60}{x}=40\]

hero (hero):

60 = 40x 6 = 4x 6/4 = x 2/3 = x

hero (hero):

in inches, not feet

OpenStudy (anonymous):

60/x = 40 x = 40/60 x = 4/6 x = 2/3 x = .67?

hero (hero):

Yes

OpenStudy (anonymous):

.7 inches. Yay!! thank you!

hero (hero):

wait, lol

OpenStudy (anonymous):

uh oh.......lol

hero (hero):

I think I did that backwards

hero (hero):

3/2 = x

hero (hero):

60 = 40x 6 = 4x 6/4 = x 3/2 = x

OpenStudy (anonymous):

1.5??

hero (hero):

Yes

OpenStudy (anonymous):

thanks!

hero (hero):

Funny thing is, even if you don't convert, you still get 1.5 I find that very strange

OpenStudy (anonymous):

ha

hero (hero):

lol